# Mamba-in-Mamba: Centralized Mamba-Cross-Scan in Tokenized Mamba Model for Hyperspectral Image Classification

### v3

@@ IV-E4 Exploring the Feature Fusion Strategy in T-Mamba
Figure 17: The comparison between using different sequential networks (i.e., RNN, GRU, LSTM, Transformer encoder, RT, Mamba, and T-Mamba). We have to note that all networks are not sharing parameters for four types of MCS in MiM model. Experimental results are listed in (a) IP dataset, (b) PU dataset, (c) HU2013 dataset, and (d) HongHu dataset. Meanwhile, the (e) shows the training time, OA(%) and model size by using those networks.
IV-E5 Comparison with Other Sequential Networks
Following the development of the T-Mamba, we conducted comparative experiments against other established sequential networks: RNN [^10] https://pytorch.org/docs/stable/generated/torch.nn.RNN.html, GRU [^11] https://pytorch.org/docs/stable/generated/torch.nn.GRU.html, LSTM [^12] https://pytorch.org/docs/stable/generated/torch.nn.LSTM.html, general Transformer encoder [^13] https://pytorch.org/docs/stable/generated/torch.nn.TransformerEncoder.html, RT (a hybrid of LSTM and Transformer) [45], and a general Mamba model [47]. These models all process vector inputs at each timestep, allowing us to integrate them directly into our system for a straightforward comparison. We have to note that all networks are not sharing parameters for four types of MCS in MiM model. All tests were performed under identical conditions, with the results shown in Fig. 17.
@@ IV-E5 Comparison with Other Sequential Networks
TABLE XII: DIFFERENT CLASSIFICATION METHODS IN TERMS OF OA ON THE INDIAN PINES DATASET WITH DATA AUGMENTATION. THE BEST ONE IS HIGHLIGHTED AS RED. THE SECOND BEST IS HIGHLIGHTED AS BLUE.
Num. Train vanilla ViT[18]     SpeFormer[19]     MAEST[21]     HiT[22]     SSFTT[26]     3DViT[27]     HUSST[28]     ViM[52]     MiM
@@
IV-E6 Further Discussion on Mamba and Transformer under Data Augmentation
@@
The results, detailed in Table XII, indicate that increasing the training sample size to nearly double improves the performance of all models, including our Mamba-based model, although the accuracy gap between methods narrows considerably. This is further evidenced by the shrinking difference in performance between SpeFormer and ViM, two purely model-driven approaches.

### v4

@@ 4.5.4 Exploring the Feature Fusion Strategy in T-Mamba
Figure 17: The comparison between using different sequential networks (i.e., RNN, GRU, LSTM, Transformer encoder, RT, Mamba, and T-Mamba). We have to note that all networks are not sharing parameters for four types of MCS in MiM model. Experimental results are listed in (a) IP dataset, (b) PU dataset, (c) HU2013 dataset, and (d) HongHu dataset.
Figure 18: The comparison on the training time, OA(%) and model size by using different sequential networks, aligning with Fig. 17.
4.5.5 Comparison with Other Sequential Networks
Following the development of the T-Mamba, we conducted comparative experiments against other established sequential networks: RNN, GRU, LSTM, general Transformer encoder, RT (a hybrid of LSTM and Transformer) [44], and a general Mamba model [46]. These models all process vector inputs at each timestep, allowing us to integrate them directly into our system for a straightforward comparison. We have to note that all networks are not sharing parameters for four types of MCS in MiM model. All tests were performed under identical conditions, with the results shown in Figs. 17 and 18.
@@ 4.5.5 Comparison with Other Sequential Networks
Table 12: Different classification methods in terms of OA on the Indian Pines dataset with data augmentation. The best one is highlighted as red. The second best is highlighted as blue.
Num. Train vanilla ViT SpeFormer MAEST HiT SSFTT 3DViT HUSST ViM MiM
@@
4.5.6 Further Discussion on Mamba and Transformer under Data Augmentation
@@
The results, detailed in Table 12, indicate that increasing the training sample size to nearly double improves the performance of all models, including our Mamba-based model, although the accuracy gap between methods narrows considerably. This is further evidenced by the shrinking difference in performance between SpeFormer and ViM, two purely model-driven approaches.